# Team Xiaomi EV-AD VLA: Caption-Guided Retrieval System for Cross-Modal Drone Navigation - Technical Report for IROS 2025 RoboSense Challenge Track 4

### v1

@@ Abstract
Cross-modal drone navigation remains a challenging task in robotics, requiring efficient retrieval of relevant images from large-scale databases based on natural language descriptions. The RoboSense 2025 Track 4 challenge addresses this challenge, focusing on robust, natural language-guided cross-view image retrieval across multiple platforms (drones, satellites, and ground cameras). Current baseline methods, while effective for initial retrieval, often struggle to achieve fine-grained semantic matching between text queries and visual content, especially in complex aerial scenes. To address this challenge, we propose a two-stage retrieval refinement method: Caption-Guided Retrieval System (CGRS) that enhances the baseline coarse ranking through intelligent reranking. Our method first leverages a baseline model to obtain an initial coarse ranking of the top 20 most relevant images for each query. We then use Vision-Language-Model (VLM) to generate detailed captions for these candidate images, capturing rich semantic descriptions of their visual content. These generated captions are then used in a multimodal similarity computation framework to perform fine-grained reranking of the original text query, effectively building a semantic bridge between the visual content and natural language descriptions. Our approach significantly improves upon the baseline, achieving a consistent 5% improvement across all key metrics (Recall@1, Recall@5, and Recall@10). Our approach win TOP-2 in the challenge, demonstrating the practical value of our semantic refinement strategy in real-world robotic navigation scenarios.
 Welcome to 2025 RoboSense Challenge!
@@ 1 Introduction
Cross-modal drone navigation represents a key advancement in autonomous robotics, enabling drones to understand and execute navigation instructions derived from natural language descriptions [5, 43, 17, 18, 16]. This technology has profound implications for applications ranging from disaster management and search and rescue operations to autonomous surveillance and environmental monitoring. Unlike traditional GPS navigation systems, natural language-guided drone navigation requires a deep understanding of spatial relationships, visual semantics, and cross-modal matching between textual descriptions and aerial imagery [9, 44, 41]. The fundamental challenge in this field lies in bridging the semantic gap between human-language descriptions and visual representations captured from aerial perspectives. Drone-viewed imagery exhibits unique characteristics, including significant viewpoint variations, scale differences, and complex spatial arrangements that differ significantly from ground-based imagery. These factors render traditional image retrieval methods inadequate for real-world drone navigation applications, where accuracy and reliability are crucial [37, 5].
Existing cross-modal retrieval methods for drone navigation primarily rely on end-to-end learning frameworks that directly map textual queries to visual features. The GeoText-1652 [5] baseline method proposed by Chu et al. represents the current state of the art in this area. It employs a multimodal framework that combines an image encoder (Swin Transformer), a text encoder (BERT) [7], and a cross-modal attention mechanism with spatially aware learning via hybrid spatial matching. Despite these advances, this baseline method still faces inherent limitations in terms of semantic precision and retrieval accuracy. Direct mapping from textual descriptions to visual embeddings often fails to capture subtle semantic relationships, especially when multiple visually similar objects are present in aerial scenes.
@@
Our experimental results on the RoboSense 2025 Track 4 challenge dataset [5] demonstrate the effectiveness of this approach, achieving a consistent 5% improvement across all key metrics (Recall@1, Recall@5, and Recall@10) compared to the baseline method. These improvements ultimately helped us secure a top-2 finish in the competition, validating the practical value of our semantic refinement strategy in real-world cross-modal drone navigation applications. The main contributions are as follows:
@@
  • Compared to baseline methods, we achieve a consistent 5% improvement on all key metrics (Recall@1, Recall@5, and Recall@10) and win TOP-2 in Cross-Modal Drone Navigation Track of the IROS 2025 RoboSense Challenge.
@@ 2.1 Cross-view Geolocalization
The task of cross-view geolocalization aims to connect visual data captured from distinct viewpoints, such as ground, aerial, or drone perspectives, to their real-world geographic positions. A central difficulty lies in designing representations that remain stable under drastic viewpoint shifts. Early studies attempted to address this by introducing strategies like partitioning images into local regions [35] or adding attention modules that highlight key landmarks [28]. More recent work explores transformer-based designs [6, 38] and hybrid models that integrate local and global features [32]. In parallel, several approaches incorporate auxiliary cues, such as pose estimation [33] or orientation information [20], while others utilize advanced augmentation or contrastive frameworks [8, 47, 50] to improve robustness. Some works also emphasize domain-specific architectures, e.g., cross-drone transformers [2]. In contrast to these directions, our study focuses on new language-driven drone navigation tasks, which naturally bridge human intention and autonomous flight by directly interpreting textual commands.
@@ 2.2 Multi-modality Alignment
Aligning text and visual information has been widely studied in vision-language retrieval [14, 19, 1, 11, 15, 17, 18, 16]. Early work emphasized architectural designs such as dual-path frameworks [49, 42, 45, 34, 48, 46, 29, 23, 4]. Later, research shifted to fine-grained alignment: adaptive gating suppressed irrelevant matches [36], and graph-based reasoning captured semantic relations [26]. More recent studies highlight robustness and relational modeling, including uncertainty-aware alignment for cross-domain video-text retrieval [10], multimodal co-attention for video-audio retrieval [13], and attentive relational aggregation for video-text retrieval [12]. With the rise of large-scale pretraining, word-region alignment [3], object-tag anchoring [27], and attribute-focused keywords [39] became common practices. At the same time, contrastive pretraining frameworks such as CLIP [31], BLIP [25], and follow-up works [22, 24, 40] achieved remarkable advances in bridging modalities. Unlike these approaches, which largely focus on semantic or global correspondence, our work emphasizes spatial-aware alignment, explicitly modeling fine-grained region–text relations. This design better suits navigation scenarios, where relative positions and local spatial cues are essential for precise drone control.
@@ 3.3 Caption-Guided Retrieval System (CGRS)
Coarse-grained Model Our coarse-grained retrieval stage is built on the GeoText-1652 baseline model [5], which adopts a multimodal framework consisting of an image encoder, a text encoder, and a cross-modal fusion module for spatially aware matching.
@@ 4.1 Benchmark
We conduct experiments on the GeoText-1652 dataset [5], a large-scale multimodal benchmark designed for natural language-guided drone geolocalization. The training set contains 50,218 images from 33 universities, with 113,562 global descriptions and 113,367 region-level bounding box-text pairs; the test set contains 54,227 images from 39 non-overlapping universities, with 154,065 global descriptions and 140,179 bounding box-text pairs. Each image is annotated with an average of 3 global descriptions (70.23 words each) and 2.62 region-level descriptions (21.6 words each), which explicitly capture spatial relationships such as “left”, “top right”, and “center”. This fine-grained spatial annotation makes GeoText-1652 [5] particularly challenging and suitable for evaluating cross-modal retrieval methods that require precise spatial understanding.
@@ 4.2 Evaluation Metrics
Following the standard protocol established by the benchmark and challenge [5], we evaluated our approach on the multimodal drone navigation task, which employs text-to-image retrieval to retrieve relevant drone-viewed images using natural language queries. We used Recall@K (R@K) as the primary evaluation metric, measuring the percentage of queries for which at least one correct image appears in the top K retrieval results. Specifically, we report R@1, R@5, and R@10 to assess the accuracy of fine-grained matching and the robustness of the overall retrieval. Higher Recall@K indicates better retrieval performance, while R@1 reflects the model’s ability to prioritize correct images.
@@ 4.3 Implementation Details
We replicated the GeoText-1652 baseline model [5] on 8 NVIDIA A800 GPUs for 5 epochs, taking approximately 30 GPU hours. We used the AdamW [30] optimizer with a learning rate of $3\times 10^{-5}$ and a spatial matching weight $\lambda=0.1$. After training, we used this baseline model to extract the top 20 candidate images for each test query. In the fine-grained reranking stage, we used GPT-4o [21] with a maximum token count of 256 to generate detailed captions for all candidate images. The average length of the generated captions was 120-150 words, capturing spatial layout and landmark details. We then used BERT [7] as the sentence encoder $\mathcal{E}_{\text{sent}}$ to compute the semantic similarity between the original query and the generated captions. After validation experiments, the hybrid fusion weight was empirically set to $\alpha=0.3$, prioritizing semantic matching based on captions while preserving the visual signals from the coarse model. The entire caption generation process takes approximately 12 hours of offline time, while the final reranking process adds almost no inference overhead (¡ 10 milliseconds per query).

### v2

# Caption-Guided Retrieval System for Cross-Modal Drone Navigation

@@ Abstract
Cross-modal drone navigation remains a challenging task in robotics, requiring efficient retrieval of relevant images from large-scale databases based on natural language descriptions. The RoboSense 2025 Track 4 challenge addresses this challenge, focusing on robust, natural language-guided cross-view image retrieval across multiple platforms (drones, satellites, and ground cameras). Current baseline methods, while effective for initial retrieval, often struggle to achieve fine-grained semantic matching between text queries and visual content, especially in complex aerial scenes. To address this challenge, we propose a two-stage retrieval refinement method: Caption-Guided Retrieval System (CGRS) that enhances the baseline coarse ranking through intelligent reranking. Our method first leverages a baseline model to obtain an initial coarse ranking of the top 20 most relevant images for each query. We then use Vision-Language-Model (VLM) to generate detailed captions for these candidate images, capturing rich semantic descriptions of their visual content. These generated captions are then used in a multimodal similarity computation framework to perform fine-grained reranking of the original text query, effectively building a semantic bridge between the visual content and natural language descriptions. Our approach significantly improves upon the baseline, achieving a consistent 5% improvement across all key metrics (Recall@1, Recall@5, and Recall@10). Our approach win TOP-3 in the challenge, demonstrating the practical value of our semantic refinement strategy in real-world robotic navigation scenarios.
@@ 1 Introduction
Cross-modal drone navigation represents a key advancement in autonomous robotics, enabling drones to understand and execute navigation instructions derived from natural language descriptions [geotext, zhang2025mapnav, hao2024mbfusion, hao2025mapfusion, hao2024your]. This technology has profound implications for applications ranging from disaster management and search and rescue operations to autonomous surveillance and environmental monitoring [geotext].
Unlike traditional GPS navigation systems, natural language-guided drone navigation requires a deep understanding of spatial relationships, visual semantics, and cross-modal matching between textual descriptions and aerial imagery [gong2025stairway, zhang2025multi, zhang2024trihelper]. The fundamental challenge in this field lies in bridging the semantic gap between human-language descriptions and visual representations captured from aerial perspectives. Drone-viewed imagery exhibits unique characteristics, including significant viewpoint variations, scale differences, and complex spatial arrangements that differ significantly from ground-based imagery [liang2025pi3det, li2025_3eed, chaney2023m3ed, kong2025eventfly, li2025seeground, bian2025dynamiccity]. These factors render traditional image retrieval methods inadequate for real-world drone navigation applications, where accuracy and reliability are crucial [wu2025evaluating, geotext, hao2025safemap, hao2025msc-bench].
Existing cross-modal retrieval methods for drone navigation primarily rely on end-to-end learning frameworks that directly map textual queries to visual features. The GeoText-1652 [geotext] baseline method proposed by Chu et al. represents the current state of the art in this area. It employs a multimodal framework that combines an image encoder (Swin Transformer), a text encoder (BERT) [bert], and a cross-modal attention mechanism with spatially aware learning via hybrid spatial matching. Despite these advances, this baseline method still faces inherent limitations in terms of semantic precision and retrieval accuracy. Direct mapping from textual descriptions to visual embeddings often fails to capture subtle semantic relationships, especially when multiple visually similar objects are present in aerial scenes.
@@
Our experimental results on the RoboSense 2025 Track 4 challenge dataset [geotext] demonstrate the effectiveness of this approach, achieving a consistent 5% improvement across all key metrics (Recall@1, Recall@5, and Recall@10) compared to the baseline method. These improvements ultimately helped us secure a top-2 finish in the competition, validating the practical value of our semantic refinement strategy in real-world cross-modal drone navigation applications. The main contributions are as follows:
@@
  • Compared to baseline methods, we achieve a consistent 5% improvement on all key metrics (Recall@1, Recall@5, and Recall@10) and win TOP-3 in Cross-Modal Drone Navigation Track of the IROS 2025 RoboSense Challenge.
@@ 2.1 Cross-view Geolocalization
The task of cross-view geolocalization aims to connect visual data captured from distinct viewpoints, such as ground, aerial, or drone perspectives, to their real-world geographic positions. A central difficulty lies in designing representations that remain stable under drastic viewpoint shifts. Early studies attempted to address this by introducing strategies like partitioning images into local regions [wang2021each] or adding attention modules that highlight key landmarks [lin2022joint]. More recent work explores transformer-based designs [dai2021transformer, yang2021cross] and hybrid models that integrate local and global features [rodrigues2022global]. In parallel, several approaches incorporate auxiliary cues, such as pose estimation [shi2022beyond] or orientation information [hu2022beyond], while others utilize advanced augmentation or contrastive frameworks [dhakal2024sat2cap, zhang2023cross, zhu2022transgeo] to improve robustness. Some works also emphasize domain-specific architectures, e.g., cross-drone transformers [chen2023cross]. In contrast to these directions, our study focuses on new language-driven drone navigation tasks, which naturally bridge human intention and autonomous flight by directly interpreting textual commands.
@@ 2.2 Multi-Modality Alignment
Aligning text and visual information has been widely studied in vision-language retrieval [hao2023dual, hao2025dada++, albanie2020end, hao2021multi, hao2023mixgen, hao2024mbfusion, hao2025mapfusion, hao2024your, chen2023clip2Scene, chen2023towards, liu2024multi, liu2023uniseg]. Early work emphasized architectural designs such as dual-path frameworks [zheng2020dual, zhang2025nava, zhang2025humanoidpano, team2025robobrain, zheng2025railway, zhang2025video, liu2025toponav, li2025vquala, cheng2025exploring]. Later, research shifted to fine-grained alignment: adaptive gating suppressed irrelevant matches [wang2019camp], and graph-based reasoning captured semantic relations [li2019visual]. More recent studies highlight robustness and relational modeling, including uncertainty-aware alignment for cross-domain video-text retrieval [hao2023uncertainty], multimodal co-attention for video-audio retrieval [hao2022listen], and attentive relational aggregation for video-text retrieval [hao2021matters]. With the rise of large-scale pretraining, word-region alignment [chen2020uniter], object-tag anchoring [li2020oscar], and attribute-focused keywords [yang2023towards] became common practices. At the same time, contrastive pretraining frameworks such as CLIP [radford2021learning], BLIP [li2022blip], and follow-up works [jia2021scaling, li2021align, zeng2021multi, sautier2022slidr, liu2023segment, kong2025largead, xu2025lima] achieved remarkable advances in bridging modalities. Unlike these approaches, which largely focus on semantic or global correspondence, our work emphasizes spatial-aware alignment, explicitly modeling fine-grained region-text relations. This design better suits navigation scenarios, where relative positions and local spatial cues are essential for precise drone control.
@@ 3.3 Caption-Guided Retrieval System (CGRS)
Coarse-grained Model Our coarse-grained retrieval stage is built on the GeoText-1652 baseline model [geotext], which adopts a multimodal framework consisting of an image encoder, a text encoder, and a cross-modal fusion module for spatially aware matching.
@@ 4.1 Benchmark
We use the official data provided by the RoboSense Challenge 2025 [robosense_challenge_2025] held at IROS 2025. This competition builds upon the legacy of the RoboDepth Challenge 2023 [robodepth_challenge_2023, kong2023robodepth] at ICRA 2023 and the RoboDrive Challenge 2024 [robodrive_challenge_2024, xie2025robobev] at ICRA 2024, continuing the collective effort to advance robust and scalable robot perception. Each track in this competition is grounded on an established benchmark designed for evaluating real-world robustness and generalization [xie2025drivebench, gong2025falcon, li2024place3d, geotext, liang2025pi3det, kong2023robo3d]. Specifically, this task is built upon the GeoText-1652 dataset [geotext] in Track 4, which benchmarks cross-modal image-text retrieval for language-guided drone navigation across drastically different viewpoints and real-world sensing conditions.
The GeoText-1652 dataset [geotext] is a large-scale multimodal benchmark designed for natural language-guided drone geolocalization. The training set contains 50,218 images from 33 universities, with 113,562 global descriptions and 113,367 region-level bounding box-text pairs; the test set contains 54,227 images from 39 non-overlapping universities, with 154,065 global descriptions and 140,179 bounding box-text pairs. Each image is annotated with an average of 3 global descriptions (70.23 words each) and 2.62 region-level descriptions (21.6 words each), which explicitly capture spatial relationships such as “left”, “top right”, and “center”. This fine-grained spatial annotation makes GeoText-1652 [geotext] particularly challenging and suitable for evaluating cross-modal retrieval methods that require precise spatial understanding.
@@ 4.2 Evaluation Metrics
Following the standard protocol established by the benchmark and challenge [geotext], we evaluated our approach on the multimodal drone navigation task, which employs text-to-image retrieval to retrieve relevant drone-viewed images using natural language queries. We used Recall@K (R@K) as the primary evaluation metric, measuring the percentage of queries for which at least one correct image appears in the top K retrieval results. Specifically, we report R@1, R@5, and R@10 to assess the accuracy of fine-grained matching and the robustness of the overall retrieval. Higher Recall@K indicates better retrieval performance, while R@1 reflects the model’s ability to prioritize correct images.
@@ 4.3 Implementation Details
We replicated the GeoText-1652 baseline model [geotext] on 8 NVIDIA A800 GPUs for 5 epochs, taking approximately 30 GPU hours. We used the AdamW [loshchilov2017decoupled] optimizer with a learning rate of $3\times 10^{-5}$ and a spatial matching weight $\lambda=0.1$. After training, we used this baseline model to extract the top 20 candidate images for each test query. In the fine-grained reranking stage, we used GPT-4o [hurst2024gpt] with a maximum token count of 256 to generate detailed captions for all candidate images. The average length of the generated captions was 120-150 words, capturing spatial layout and landmark details. We then used BERT [bert] as the sentence encoder $\mathcal{E}_{\text{sent}}$ to compute the semantic similarity between the original query and the generated captions. After validation experiments, the hybrid fusion weight was empirically set to $\alpha=0.3$, prioritizing semantic matching based on captions while preserving the visual signals from the coarse model. The entire caption generation process takes approximately 12 hours of offline time, while the final reranking process adds almost no inference overhead (¡ 10 milliseconds per query).